OpenStudy (anonymous):

The conical hat shown here is a traditional sun-protective straw hat originating in eastern Asia. What is the approximate outer surface area of one of these hats that has a radius of 9 inches and a slant height of twelve inches? Keep in mind that there is an opening in the hat where the base surface would normally be. Show the steps of your solution and explain your work. woman wearing a straw hat that has the shape of a wide cone.

OpenStudy (lxelle):

The surface area of a cone (excluding the base) is pi*r*s, where r = radius and s = slant height We have pi*9*12 = 96*pi =???

OpenStudy (dangerousjesse):

Be sure to give credit before you give someone else's answer :)

OpenStudy (anonymous):

Thanks @LXelle I am starting to get your methods! :)

OpenStudy (lxelle):

:))